22) At a youth camp, the counsellor sent a frog-catching team to the swamp to catch...

22)

At a youth camp, the counsellor sent a frog-catching team to the swamp to catch frogs; he assumed that the team was equally likely to catch male as female frogs. But of the 14 frogs that they brought back, 5 were females. What is the probability of this?

Select one:

( ) a. 0.24

( ) b. 0.36

( ) c. 0.06

( ) d. 0.12
Answer #1

Answer)

As there are fixed number of trials and probability of each and every trial is same and independent of each other

Here we need to use the binomial formula

P(r) = ncr*(p^r)*(1-p)^n-r

Ncr = n!/(r!*(n-r)!)

N! = N*n-1*n-2*n-3*n-4*n-5........till 1

For example 5! = 5*4*3*2*1

Special case is 0! = 1

P = probability of single trial = 0.5

N = number of trials = 14

R = desired success = 5

P(5) = 14c5*(0.5^5)*(1-0.5)^14-5 = 0.12